# Hierarchical Empowerment: Towards Tractable Empowerment-Based Skill-Learning

### v1

@@ 3 Hierarchical Empowerment
We introduce a framework, Hierarchical Empowerment, that makes two improvements to using goal-conditioned RL as a variational lower bound on mutual information and empowerment. The first improvement we make is a new variational lower bound, Goal-Conditioned Empowerment, that uses the same fixed variational distribution as Choi et al. [2021], but also learns the distribution of skills. The second improvement is to integrate a hierarchical architecture that makes computing the empowerment over long time horizons more tractable.
@@ 3.1 Goal-Conditioned Empowerment
$\mathcal{E}^{GCE}(s_{0},\theta,\phi)=\max_{\pi_{\theta},p_{\phi}}I^{GCE}(Z;S_{n}|s_{0}).$ (5)
@@
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathrm{E}_{z\sim p_{\phi}(z|s_{0}),s_{n}\sim p_{\theta}(s_{n}|s_{0},z)}[\log q(z|s_{0},s_{n})].$ (6)
Next, we will analyze the variational lower bound objective in equation 6 with respect to the goal-space policy and goal-conditioned policy separately.
@@
An immediate challenge to optimizing Equation 6 with respect to the goal space parameters $\phi$ is that $\phi$ occurs in the expectation distribution. The reparameterization trick [Kingma and Welling, 2014] provides a solution to this issue by transforming the original expectation into an expectation with respect to exogenous noise. In order to use the reparameterization trick, the distribution of goal states needs to be a location-scale probability distribution (e.g., the uniform or the normal distributions). For the remainder of the paper, we will assume $p_{\phi}(z|s_{0})$ is a uniform distribution.
Applying the reparameterization trick to the expectation term in Equation 6 and inserting the fixed variance Gaussian variational distribution used by [Choi et al., 2021], the variational lower bound objective becomes
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathrm{E}_{\epsilon\sim\mathcal{U}^{d}}[\mathrm{E}_{s_{n}\sim p(s_{n}|s_{0},z)}[\log\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})]],$ (7)
@@
in which $d$ is the dimensionality of the goal space and $z=g(\epsilon,\mu_{\phi}(s_{0}))$ is reparameterized to be a function of the exogenous unit uniform random variable $\epsilon$ and $\mu_{\phi}(s_{0})$, which outputs a vector containing the location-scale parameters of the uniform goal-space distribution. Specifically, for each dimension of the goal space, $\mu_{\phi}(s_{0})$ will output (i) the location of the center of the uniform distribution and (ii) the length of the half width of the uniform distribution. Note also that in our implementation, instead of sampling the skill $z$ from the fixed variational distribution (i.e., $\mathcal{N}(z;s_{n},\sigma_{0})$), we sample the sum $z+s_{0}$ (i.e., $\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})$). This small change will encourage the goal $z$ to reflect the desired change from the starting state $s_{0}$. The mutual information objective in Equation 7 can be further simplified to
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathrm{E}_{\epsilon\sim\mathcal{U}^{d}}[R(s_{0},\epsilon,\mu_{\phi}(s_{0}))],$ (8)
$R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ $=\mathrm{E}_{s_{n}\sim p(s_{n}|s_{0},\epsilon,\mu_{\phi}(s_{0}))}[\log\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})].$
@@
Equation 8 has the same form as a maximum entropy bandit problem. The entropy term $H_{\phi}(Z|s_{0})$ encourages the goal space policy $\mu_{\phi}$ to output larger goal spaces. The reward function $\mathrm{E}_{\epsilon\sim\mathcal{U}^{d}}[R(s_{0},\epsilon,\mu_{\phi}(s_{0}))]$ in the maximum entropy bandit problem encourages the goal space to contain achievable goals. This is true because $R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ measures how effective the goal-conditioned policy is at achieving the particular goal $z$, which per the reparameterization trick is $g(\epsilon,\mu_{\phi}(s_{0}))$. The outer expectation $\mathrm{E}_{\epsilon\sim\mathcal{U}^{d}}[\cdot]$ then averages $R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ with respect to all goals in the learned goal space. Thus, goal spaces that contain more achievable goals will produce higher rewards.
@@
The goal space policy $\mu_{\phi}$ can be optimized with Soft Actor-Critic [Haarnoja et al., 2018], in which a critic $R_{\omega}(s_{0},\epsilon,a)$ is used to approximate $R(s_{0},\epsilon,a)$ for which $a$ represents actions by the goal space policy. In our implementation, we fold in the entropy term into the reward, and estimate the expanded reward with a critic. We then use Stochastic Value Gradient [Heess et al., 2015] to optimize.
Under this optimization strategy, the objective for training the critic is
$J(\omega)$ $=\mathrm{E}_{(s_{0},\epsilon,a)\sim\beta}[(R_{\omega}(s_{0},\epsilon,a)-\text{Target})^{2}],$ (9)
Target $=\mathrm{E}_{s_{n}\sim p(s_{n}|s_{0},z)}[\log\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})-\log p(z|s_{0})].$
In Equation 9, the learned critic is trained to be closer to a target value using $(s_{0},\epsilon,a)$ tuples from a buffer $\beta$. The target is obtained by first executing the skill $z=g(\epsilon,a)$, in which $a$ is a goal space action (i.e., the vector of location-scale parameters for the uniform distribution goal space), and then sampling the terminal state $s_{n}$. The target is then the sum of (i) the $\log\mathcal{N}(s_{0}+z;s_{n},\sigma_{0})$, which measures how effective the goal-conditioned policy is at achieving the goal state $s_{0}+z$ and (ii) $-\log p(z|s_{0})$, which comes from the entropy reward that was folded in.
The gradient for the goal-space policy $\mu_{\phi}$ can then be determined by passing the gradient through the critic using the chain rule:
$\nabla_{\phi}I^{GCE}(Z;S_{n}|s_{0})=\nabla_{\phi}\mu_{\phi}(s_{0})\mathrm{E}_{\epsilon\sim\mathcal{U}^{d}}[\nabla_{a}R_{\omega}(s_{0},\epsilon,a)|_{a=\mu_{\phi}(s_{0})}].$ (10)
@@
With respect to the goal-conditioned policy $\pi_{\theta}$, the mutual information objective reduces to
$\mathrm{E}_{z\sim p_{\phi}(z|s_{0}),s_{n}\sim p_{\theta}(s_{n}|s_{0},z)}[\log q(z|s_{0},s_{n})],$ (11)
because the entropy term can be ignored. This is same objective as the goal-conditioned RL variational lower bound introduced by Choi et al. [2021] except the outer expectation is with respect to a learned goal space. We optimize this objective with respect to the deterministic goal-conditioned policy with the Deterministic Policy Gradient [Silver et al., 2014].
@@
In our implementation, we train the Goal-Conditioned Empowerment mutual information objective by iterating between updates to the goal-conditioned policy and the goal-space policy—similar to an automated curriculum of goal-conditioned RL problems. In each step of the curriculum, the goal-conditioned policy is trained to reach goals within and nearby the current goal space. The goal space is then updated to reflect the current reach of the goal-conditioned policy. Detailed algorithms for how the goal-conditioned actor-critic and the goal-space actor-critic are updated are provided in section D of the Appendix.
@@ 3.2 Hierarchical Architecture
General purpose agents will need to have both short and long time horizon skills. One way to learn both short and temporally extended skills is to be able to compute the long-term empowerment of a state (i.e., learn skills that target the state space that is reachable over long horizons). A key problem with Goal-Conditioned Empowerment is that it is not a practical solution for computing long-term empowerment because it is difficult to optimize temporally extended policies with RL. However, Goal-Conditioned Empowerment can be effective at learning skills containing short sequences of actions. We introduce a solutions that takes advantage of this capability.
To scale Goal-Conditioned Empowerment to longer horizons we borrow ideas from Goal-Conditioned Hierarchical Reinforcement Learning (GCHRL), sometimes referred to as Feudal RL [Dayan and Hinton, 1992, Levy et al., 2017, 2019, Nachum et al., 2018, Zhang et al., 2021]. GCHRL approaches have shown that temporally extended policies can be learned by nesting multiple short sequence policies. In the Hierarchical Empowerment framework, this approach is implemented as follows. First, the designer implements $k$ levels of Goal-Conditioned Empowerment, in which $k$ is a hyperparameter set by the designer. That is, the agent will learn $k$ goal-conditioned and goal space policies $(\pi_{\theta_{0}},\mu_{\phi_{0}},\dots,\pi_{\theta_{k-1}},\mu_{\phi_{k-1}})$. Second, for all levels above the bottom level, the action space is set to the learned goal space from the level below. (The action space for the bottom level remains the primitive action space.) For instance, for an agent with $k=2$ levels of skills, the action space for the top level goal-conditioned policy $\pi_{\theta_{1}}$ at state $s$ will be set to the goal space output by $\mu_{\phi_{0}}(s)$. The top level goal-space $\mu_{\phi_{1}}$ and goal-conditioned policies $\pi_{\theta_{1}}$ will thus seek to learn the largest space of states reachable in $n$ subgoals, in which each subgoal can contain at most $n$ primitive actions. With this nested structure inspired by GCHRL, the time horizon of the computed empowerment can grow exponentially with $k$, and no goal-conditioned policy is required to learn a long sequence of actions, making optimizing Goal-Conditioned Empowerment with RL more tractable. Figure 2 visualizes the proposed hierarchical architecture in one of our domains.
@@ 3.3 Limitations
The Hierarchical Empowerment framework has two significant limitations as currently constructed. The first is that it can only be applied to domains in which the state space has large contiguous regions of reachable states that can be covered by $d$-dimensional goal space boxes. In the framework, the uniform goal space distribution will only expand to include achievable states. If regions adjacent to the goal space are not reachable, the goal space will not expand. Consequently, if the state space contains large regions of unreachable states (e.g., a pixel space), the framework will not be able to learn a meaningful space of skills.
The second limitation is that a model of the transition dynamics is required. The most important need for the model is to simulate the temporally extended subgoal actions so that the higher level goal-conditioned policies can be trained with RL. We also take advantage of the model to run episodes in parallel to make training more efficient. However, the existing empowerment-based skill-learning approaches that learn the variational distribution $q_{\psi}(z|s_{0},s_{n})$ also require a model in large domains to obtain unbiased gradients of the maximum likelihood objective. The maximum likelihood objective requires sampling large batches of skills from large sets of skill starting states and then executing the current skill-conditioned policy. In large domains, a model of the transition dynamics is required for this to be feasible.
@@ 4 Experiments
The purpose of our experiments is to evaluate the two claims of our framework. The first claim is that Goal-Conditioned Empowerment, which learns skills using an automated curriculum of goal-conditioned RL problems, can more effectively compute empowerment (i.e., learn the largest space of distinct skills) than existing empowerment-based skill learning approaches that rely on a learned variational distribution.
For this evaluation, we compare Goal-Conditioned Empowerment to Diversity Is All You Need (DIAYN) [Eysenbach et al., 2019]. The skill-learning objective used by DIAYN is very similar to the typical empowerment-based skill-learning objective described in section 2.3. DIAYN employs a fixed space of skills $p(z)$ and uses a learned variational distribution $q_{\psi}(z|s_{t})$ trained with maximum likelihood. One difference from the prototypical empowerment objective is that DIAYN optimizes a regularized version of that objective in which the reward $\log q_{\psi}(z|s_{t})$ occurs at each time step. DIAYN also includes the entropy term $H(\pi_{\theta}(a_{t}|s_{t},z))$ to help skill-conditioned policy explore.
We make some changes to DIAYN to make the comparison more fair. One change is we implement DIAYN with a continuous skill space in which $p(z)$ is sampled from a uniform distribution. Also, as in Goal-Conditioned Empowerment, we condition the skills on a start state (i.e., $p(z)\rightarrow p(z|s_{0})\text{ and }q_{\phi}(z|s_{t})\rightarrow q_{\phi}(z|s_{0},s_{t})$) and limit the number of primitive actions included in the skill. With these changes, the implementation of DIAYN is similar to HIDIO [Zhang et al., 2021], except HIDIO conditions the learned variational distribution on entire trajectories and not just individual states.
The second claim of our framework that our experiments seek to evaluate is that hierarchy is required to compute long-term empowerment. To assess this claim, we compare Hierarchical Empowerment agents that learn one, two, and three levels of skills.
@@ 4.1 Domains
Although there are some emerging benchmarks for unsupervised skill-learning [Laskin et al., 2021, Gu et al., 2021, Kim et al., 2021], we cannot evaluate Hierarchical Empowerment in these benchmarks due to the limitations of the framework. The benchmarks either (i) do not provide access to a model of the transition dynamics, (ii) use downstream tasks that require achieving goals states with too many dimensions, or (iii) do not provide downstream goal-reaching tasks.
Instead, we implement our own domains in the popular Ant and Half Cheetah environments. We implement the typical open field Ant and Half Cheetah domains involving no barriers. We also implement environments with barriers. In our Ant Cage and Half Cheetah Cage domains, the agent must press a button to remove a barrier. The purpose of implementing domains with barriers was to show that the learned goal spaces do adjust to the shape of the reachable state space. All our domains were implemented using Brax [Freeman et al., 2021], which provides access to a model of the transition dynamics.
Each experiment involves two phases of training. In the first phase, there is no external reward, and the agent simply optimizes its mutual information objective to learn skills. In the second phase, the agent learns a policy over the skills to solve a downstream task. In our experiments, the downstream task is to achieve a goal state uniformly sampled from a hand-designed continuous space of goals. Additional details on how the first and second phases are implemented are provided in section E of the Appendix.

### v2

@@ 3 Hierarchical Empowerment
We introduce a framework, Hierarchical Empowerment, that makes two improvements to using goal-conditioned RL as a variational lower bound on mutual information and empowerment. The first improvement we make is a new variational lower bound, Goal-Conditioned Empowerment, that uses the same fixed variational distribution as Choi et al. (2021), but also learns the distribution of skills. The second improvement is to integrate a hierarchical architecture that makes computing the empowerment over long time horizons more tractable.
@@ 3.1 Goal-Conditioned Empowerment
$\mathcal{E}^{GCE}(s_{0},\theta,\phi)=\max_{p_{\phi},\pi_{\theta}}I^{GCE}(Z;S_{n}|s_{0}).$ (4)
@@
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathbb{E}_{z\sim p_{\phi}(z|s_{0}),s_{n}\sim p_{\theta}(s_{n}|s_{0},z)}[\log q(z|s_{0},s_{n})].$ (5)
@@
Next, we will analyze the variational lower bound objective in equation 5 with respect to the goal-space policy and goal-conditioned policy separately. An immediate challenge to optimizing Equation 5 with respect to the goal space parameters $\phi$ is that $\phi$ occurs in the expectation distribution. The reparameterization trick (Kingma & Welling, 2014) provides a solution to this issue by transforming the original expectation into an expectation with respect to exogenous noise. In order to use the reparameterization trick, the distribution of goal states needs to be a location-scale probability distribution (e.g., the uniform or the normal distributions). We will assume $p_{\phi}(z|s_{0})$ is a uniform distribution for the remainder of the paper. Applying the reparameterization trick to the expectation term in Equation 5 and inserting a fixed variance Gaussian variational distribution similar to the one used by (Choi et al., 2021), the variational lower bound objective becomes
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathbb{E}_{\epsilon\sim\mathcal{U}^{d}}[\mathbb{E}_{s_{n}\sim p(s_{n}|s_{0},z)}[\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0})]],$ (6)
@@
in which $d$ is the dimensionality of the goal space and $z=g(\epsilon,\mu_{\phi}(s_{0}))$ is reparameterized to be a function of the exogenous unit uniform random variable $\epsilon$ and $\mu_{\phi}(s_{0})$, which outputs a vector containing the location-scale parameters of the uniform goal-space distribution. Specifically, for each dimension of the goal space, $\mu_{\phi}(s_{0})$ will output (i) the location of the center of the uniform distribution and (ii) the length of the half width of the uniform distribution. Note also that in our implementation, instead of sampling the skill $z$ from the fixed variational distribution (i.e., $\mathcal{N}(z;s_{n},\sigma_{0})$), we sample the sum $h(s_{0})+z$, in which $h(\cdot)$ is the function that maps the state space to the skill (i.e. goal) space. The change to sampling the sum $h(s_{0})+z$ will encourage the goal $z$ to reflect the desired change from the starting state $s_{0}$. The mutual information objective in Equation 6 can be further simplified to
$I^{GCE}_{\theta,\phi}(Z;S_{n}|s_{0})$ $=H_{\phi}(Z|s_{0})+\mathbb{E}_{\epsilon\sim\mathcal{U}^{d}}[R(s_{0},\epsilon,\mu_{\phi}(s_{0}))],$ (7)
$R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ $=\mathbb{E}_{s_{n}\sim p(s_{n}|s_{0},\epsilon,\mu_{\phi}(s_{0}))}[\log\mathcal{N}(h(s_{0})+z;s_{n},\sigma_{0})].$
@@
Equation 7 has the same form as a maximum entropy bandit problem. The entropy term $H_{\phi}(Z|s_{0})$ encourages the goal space policy $\mu_{\phi}$ to output larger goal spaces. The reward function $\mathbb{E}_{\epsilon\sim\mathcal{U}^{d}}[R(s_{0},\epsilon,\mu_{\phi}(s_{0}))]$ in the maximum entropy bandit problem encourages the goal space to contain achievable goals. This is true because $R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ measures how effective the goal-conditioned policy is at achieving the particular goal $z$, which per the reparameterization trick is $g(\epsilon,\mu_{\phi}(s_{0}))$. The outer expectation $\mathbb{E}_{\epsilon\sim\mathcal{U}^{d}}[\cdot]$ then averages $R(s_{0},\epsilon,\mu_{\phi}(s_{0}))$ with respect to all goals in the learned goal space. Thus, goal spaces that contain more achievable goals will produce higher rewards.
@@
The goal space policy $\mu_{\phi}$ can be optimized with Soft Actor-Critic (Haarnoja et al., 2018), in which a critic $R_{\omega}(s_{0},\epsilon,a)$ is used to approximate $R(s_{0},\epsilon,a)$ for which $a$ represents actions by the goal space policy. In our implementation, we fold in the entropy term into the reward, and estimate the expanded reward with a critic. We then use Stochastic Value Gradient (Heess et al., 2015) to optimize. Section F of the Appendix provides the objective for training the goal-space critic. The gradient for the goal-space policy $\mu_{\phi}$ can be determined by passing the gradient through the critic using the chain rule:
$\nabla_{\phi}I^{GCE}(Z;S_{n}|s_{0})=\nabla_{\phi}\mu_{\phi}(s_{0})\mathbb{E}_{\epsilon\sim\mathcal{U}^{d}}[\nabla_{a}R_{\omega}(s_{0},\epsilon,a)|_{a=\mu_{\phi}(s_{0})}].$ (8)
@@
With respect to the goal-conditioned policy $\pi_{\theta}$, the mutual information objective reduces to a goal-conditioned RL problem because the entropy term can be ignored. We optimize this objective with respect to the deterministic goal-conditioned policy with Deterministic Policy Gradient (Silver et al., 2014).
@@
In our implementation, we optimize the Goal-Conditioned Empowerment mutual information objective by alternating between updates to the goal-conditioned and goal-space policies—similar to an automated curriculum of goal-conditioned RL problems. In each step of the curriculum, the goal-conditioned policy is trained to reach goals within and nearby the current goal space. The goal space is then updated to reflect the current reach of the goal-conditioned policy. Detailed algorithms for how the goal-conditioned actor-critic and the goal-space actor-critic are updated are provided in section D of the Appendix.
@@ 3.2 Hierarchical Architecture
To scale Goal-Conditioned Empowerment to longer horizons we borrow ideas from Goal-Conditioned Hierarchical Reinforcement Learning (GCHRL), sometimes referred to as Feudal RL (Dayan & Hinton, 1992; Levy et al., 2017; 2019; Nachum et al., 2018; Zhang et al., 2021). GCHRL approaches have shown that temporally extended policies can be learned by nesting multiple short sequence policies. In the Hierarchical Empowerment framework, this approach is implemented as follows. First, the designer implements $k$ levels of Goal-Conditioned Empowerment, in which $k$ is a hyperparameter set by the designer. That is, the agent will learn $k$ goal-conditioned and goal space policies $(\pi_{\theta_{0}},\mu_{\phi_{0}},\dots,\pi_{\theta_{k-1}},\mu_{\phi_{k-1}})$. Second, for all levels above the bottom level, the action space is set to the learned goal space from the level below. (The action space for the bottom level remains the primitive action space.) For instance, for an agent with $k=2$ levels of skills, the action space for the top level goal-conditioned policy $\pi_{\theta_{1}}$ at state $s$ will be set to the goal space output by $\mu_{\phi_{0}}(s)$. Further, the transition function at level $i>0$, $T_{i}(s_{t+1}|s_{t},a_{i_{t}})$, fully executes subgoal action $a_{i}$ proposed by level $i$. For instance, for a $k=2$ level agent, sampling the transition distribution at the higher level $T_{1}(s_{t+1}|s_{t},a_{1_{t}})$ for subgoal action $a_{1_{t}}$ involves passing subgoal $a_{1_{t}}$ to level 0 as the level 0’s next goal. The level 0 goal-conditioned policy then has at most $n$ primitive actions to achieve goal $a_{1_{t}}$. Thus for a 2-level agent, the top level goal-space $\mu_{\phi_{1}}$ and goal-conditioned policies $\pi_{\theta_{1}}$ will thus seek to learn the largest space of states reachable in $n$ subgoals, in which each subgoal can contain at most $n$ primitive actions. With this nested structure inspired by GCHRL, the time horizon of the computed empowerment can grow exponentially with the number of levels $k$, and no goal-conditioned policy is required to learn a long sequence of actions, making optimizing Goal-Conditioned Empowerment with RL more tractable. Figure 2 visualizes the proposed hierarchical architecture in one of our domains. Figure E in the Appendix provides some additional detail on the hierarchical architecture.
@@ 3.3 Limitations
The Hierarchical Empowerment framework has two significant limitations as currently constructed. The first is that the framework can only learn long horizon skills in domains in which the state space has large regions of reachable states that can be covered by $d$-dimensional goal space boxes. This limitation results from the assumption that the learned goal space takes the form of a uniform distribution across a $d$-dimensional subset of the state space. In the framework, the uniform goal space distribution will only expand to include achievable states. If regions adjacent to the goal space are not reachable, the goal space will not expand. Consequently, if the state space contains large regions of unreachable states (e.g., a pixel space or a maze), the framework will not be able to learn a meaningful space of skills.
The second limitation is that a model of the transition dynamics is required. The most important need for the model is to simulate the temporally extended subgoal actions so that (state, subgoal action, next state, reward) tuples can be obtained to train the higher level goal-conditioned policies with RL. However, the existing empowerment-based skill-learning approaches that learn the variational distribution $q_{\psi}(z|s_{0},s_{n})$ also require a model in large domains to obtain unbiased gradients of the maximum likelihood objective. The maximum likelihood objective requires sampling large batches of skills $z$ from large sets of skill starting states $s_{0}$ and then executing the current skill-conditioned policy to obtain $s_{n}$. In large domains, a model of the transition dynamics is needed for this to be feasible.
@@ 4 Experiments
The purpose of our experiments is to evaluate the two claims of our framework. The first claim is that Goal-Conditioned Empowerment can more effectively compute empowerment (i.e., learn the largest space of distinct skills) than existing empowerment-based skill learning approaches that rely on a learned variational distribution. For this evaluation, we compare Goal-Conditioned Empowerment to DIAYN (Eysenbach et al., 2019), HIDIO (Zhang et al., 2021), and DADS (Sharma et al., 2020). The skill-learning objectives used by DIAYN and HIDIO are very similar to the typical empowerment-based skill-learning objective described in section 2.3. The key differences between DIAYN and HIDIO is the structure of the learned variational distributions. DIAYN learns the distribution $q_{\psi}(z|s_{0},s_{t})$ (i.e., each skill $z$ should target a specific state). In our implementation of HIDIO, we used the distribution that the authors note generally performed the best in their experiments: $q_{\phi}(z|s_{0},a_{t-1},s_{t})$ (i.e., each skill targets an (action, next state) combination). DADS optimizes a variational lower bound to a related objective to Empowerment: $I(Z;S_{1},S_{2},\dots,S_{n}|s_{0})$ (i.e., each skill should target a trajectory of states). We discuss this objective in more detail in section H of the Appendix including how DADS also does not provide a strong signal for skills to differentiate when skills overlap. The second claim of our framework that our experiments seek to evaluate is that hierarchy is helpful for computing long-term empowerment. To assess this claim, we compare Hierarchical Empowerment agents that learn one, two, and three levels of skills.
@@ 4.1 Domains
Although there are some emerging benchmarks for unsupervised skill learning (Laskin et al., 2021; Gu et al., 2021; Kim et al., 2021), we cannot evaluate Hierarchical Empowerment in these benchmarks due to the limitations of the framework. The benchmarks either (i) do not provide access to a model of the transition dynamics or (ii) do not contain state spaces with large contiguous reachable regions that can support the growth of $d$-dimensional goal space boxes. For instance, some frameworks involve tasks like flipping or running in which the agent needs to achieve particular combinations of positions and velocities. In these settings, there may be positions outside the current goal space the agent can achieve, but if the agent cannot achieve those positions jointly with the velocities in the current goal space, the goal space may not expand.
Instead, we implement our own domains in the popular Ant and Half Cheetah environments. We implement the typical open field Ant and Half Cheetah domains involving no barriers, but also environments with barriers. All our domains were implemented using Brax (Freeman et al., 2021b), which provides access to a model of the transition dynamics. In all domains, the learned goal space is limited to the position of the agent’s torso ((x,y) positions in Ant and (x) in Half Cheetah). When limited to just the torso position, all the domains support large learned goal spaces. Also, please note the default half cheetah gait in Brax is different than the default gait in MuJoCo (Todorov et al., 2012). We provide more detail in section J of the Appendix.
Each experiment involves two phases of training. In the first phase, there is no external reward, and the agent simply optimizes its mutual information objective to learn skills. In the second phase, the agent learns a policy over the skills to solve a downstream task. In our experiments, the downstream task is to achieve a goal state uniformly sampled from a hand-designed continuous space of goals. Additional details on how the first and second phases are implemented are provided in section G of the Appendix.